Unit 5: Gases Ato Basic Gas Relations . onsider the Ideal Gas Law: where n is the number of moles, P is the pressure in atm nRT wnere n is the number of moles, P is the pressure in atm, is the vol ume in L, T is the absolute (Kelvin) temperature, and R = 0.082 L atm/mole K ote: 1. Parameters that are on oppos1te sign are directly proportional to ea 2. Parameters that are on the same $1 2. Parameters that are on the same side of te equ inversely proportional to each othe 5-1 T r, 6 x 105つmolecules of a gas container,en-x-103 5-1 In a fixed volume container, 6 x 1ioLeeui molecules molecules of gas were volume of a gas exert a pressure of 12 N/m. If 12 x 10 molecules of added to the container, what would the pressure becom e? B. Quantitative Calculati tative Calculations with the Ideal Gas Law convert OC to K by adding 273 to the oc temp. Note: Always Must convert mmHg or to pressure by 760, and make sure volume is in convert mmHg or torr to atm by dividing the mmHg or torr ow many moles would occupy a volume of 500 mL at a pressure -2 H of 1520 mmHg and a temperature of 25°c? C. Avogadros Law Equal volumes of different gases, if measured at the same temperature and pressure, contain equal numbers of gas molecules. Hence, can do stoichiometry (ratio) calculations based on a balanced equation in volume units so long as the gases are measured at the same temperature and pressure. 5-3 For the reaçtion given below, how many Lters of ammonia iven below, how many Lfers of ammonia form when 7.5 kiters of hydrogen reacts? N2 (g) + 3H2 (g) - > 2 NH 3 (g) D. Daltons Law of Partial Pressures/ The total pressure of a gas is a simple sum of the partial pressures of the individual gases that are present in the gas mixture. 5-4 A gas mixture containg helium, oxygen, and nitrogen has a total pressure of 12 atm. If the helium has a partial pressure of 2 atm and the oxygen has a partial pressure of 3 atm, what is the partial pressure of the nitrogen gas? Note that for a gas collected over water (collected wet) one must subtract the vapor pressure of the water from the total pressure to get the pressure of the gas.


show calculations. 5-1, please.
Answer #1

Given

Initial

No. of molecules = 6 * 1023

No. of moles n1 = 6 * 1023 / 6.023 * 1023 = 0.9962 moles

Pressure P1 = 12 N/m2

Final

No. of molecules = 12 * 1023

No. of moles n2 = 12 * 1023 / 6.023 * 1023 = 1.9924 moles

PV = nRT

at constant T and V

P is proportional to n

P1 / P2 = n1/ n2

12 N/m2 /P2 = 0.9962 moles / 1.9924 moles

P2 = 6 N/m2 Answer